OpenStudy (anonymous):

how to write an equation of a line when the points are (0,-4) and (-5,-7)

OpenStudy (anonymous):

a straight line is composed of Y=mx+c where m is gradient c is y intercept

OpenStudy (anonymous):

to find the gradient it's change in y divided by change in x

OpenStudy (anonymous):

first find out the slope of the line \[\frac{ y2-y1 }{ x2-x1 }\] and use that slope (m) to find out y = mx +b, and input your points to find out the y-intercept (b) and you will be able to form a correct general formula for the line

OpenStudy (anonymous):

to find the intercept you sub in the values of x and y and your gradient in the equation y=mx+c and u get your value of c